84=3/4x+30
find x.

Answer by checkley77(12844)

84=3/4x+30
84=3X/4+30
3X.4=84-30
3X/4=54
3X=54*4
3X=216
X=216/3
X=72 ANS.
PROOF:
84=3*72/4+30
84=216/4+30
84=54+30
84=84